1 2 3 4 1 Product A Product B Product C 2 Machine I 15 15...

1

2

3

4

1

Product A

Product B

Product C

2

Machine I

15

15

10

3

Machine II

5

5

10

4

Machine III

14

14

5

5

Machine IV

10

10

5

6

Maximum sales

100

26.7

50

7

Sales price (in.$).

65

71

90

8

Material costs (in.$

20

35

30

9

Throughput.(in.$).

If the operating cost is $2500 then;  

Select one:

a. The profit (P) earned by firm is $ 5960

b. The profit (P) earned by firm is $ 5460

c. The profit (P) earned by firm is $ 7380

d. The profit (P) earned by firm is $ 5000

e. The profit (P) earned by firm is $ 0
Answer #1

Sales pricefor material -

  1. A = 65*100= 6500
  2. B= 71*26.7= 1895.7
  3. C= 90*50= 4500

Total sales price = 12895.7

Material cost -

  1. A= 20*100= 2000
  2. B =35*26.7= 934.5
  3. C = 30*50 =1500

Total Material cost = 4434.5

Total cost price = total material cost + total operating cost

                           = 4434.5 + 2500 = 6934.5

therefore,

Profit = total sales price - total cost price

          = 12895.7 - 6934.5 = 5961.2  

approx = $5960

The profit (P) earned by firm is $ 5960